# Voting-Based Semi-Parallel Proof-of-Work Protocol

### v3

@@ E-2 Analysis
(Guo-Ren[31]) Under AHPODH, if any attack succeeds in violating a transaction’s safety then the private mining attack also succeeds in violating that transaction’s safety.
@@
Let $\lambda_{B}=\frac{1}{T_{B}}$ be the block generation rate where $T_{B}$ denotes the block interarrival time. Let $\lambda_{P}$ denote the proof generation rate. Let $\beta$ denote the fraction of honest hashpower in the system. As honest miners need $L$ sequential proofs in the rigged model, we simply assume
@@
Let, $S_{L}\sim NegBin(L,\beta)$ denote the sum of $L$ i.i.d. geometric random variables with success probability $\beta$. Let $S_{\delta_{L}}\sim\sum_{i=1}^{L}Poisson(\lambda_{P}\delta)$ denote the sum of $L$ i.i.d. Poisson random variables with rate $\lambda_{P}\delta$, whereas $S_{\Delta_{2}}\sim\sum_{i=1}^{2}Poisson(\lambda_{P}\Delta)$.
@@
Again, for simplicity, we assume the same rigged model of [31], even though it is not possible for the adversary to mount a balanced attack after $\tau_{c}$ anymore [41]. Making the same mild assumption of ignoring the ledger delays after block $h+2$ for simplicity and incorporating all the proofs to be mined after $\tau_{c}$ into a single proof tree, the number of proofs that the adversary can make up for its deficit is denoted as $M_{deficit}\sim M_{lead}$, see [31, 32, 42, 41].
@@ Appendix F MDP Models for the Proposed Protocol
In this appendix, we assume the reader is familiar with the $\epsilon$-optimal selfish mining strategies of [5]. We avoid presenting the same average reward ratio MDP model and the results. We simply refer the reader to check relevant parts of the paper [5, Section 2, Section 4.1].
@@
We consider the following modified objective function representing the relative payoff, i.e.,
@@
F-A Simulations
To evaluate the performance of the reward split MDP and compare it with the Nakamoto consensus MDP of [5], we run the MDP models for $\gamma=0,0.2,0.4,0.5,0.6,0.8$ and $\alpha$ values range from $0.02$ to $0.48$ with $0.02$ spacing between them with $10^{-5}$ precision. It is clear that, reward splitting helps adversary when $\gamma$ is low as matching always gives a half reward when $(a,h,fork)=(1,1,relevant)$, whereas the adversarial relative reward in Nakamoto consensus increases with increasing $\gamma$ since matching succeeds more often. Thus, even when $\gamma=0.8$, the profitability threshold in our protocol stays around $\alpha\approx 0.28$.
@@ F-A Simulations
Figure 10: Reward split MDP vs Nakamoto MDP. [5]
@@ Appendix G Transaction Split Rule
It is in the interest of miner $M$ to create an initiator proof referring to its own proposed ledger instead of the ledger offering the most fees if,
@@
    If it is successful to mine an initiator referring to the ledger offering $F_{M}$ fees before others, which happens with probability $\alpha$, it shares the initiator immediately. Other miners accept the initiator of $M$ and mine on top of it since it is the chain with heaviest work.
@@
For the rest of the L1L-1 incremental proofs after the initiator, the miner $M$, creates $(L-1)\alpha$ more proofs on average. The expected transaction fees rewards is,
$F_{M}\cdot\alpha\left(r+\frac{1-r}{L}+\frac{1-r}{L}(L-1)\alpha\right)+$
$F_{M^{\prime}}\cdot(1-\alpha)\left(\frac{1-r}{L}(L-1)\alpha\right).$ (27)
On the other hand if it simply tries to mine an initiator proof referring to the ledger offering $F_{M^{\prime}}$ fees, the expected transaction fees rewards is,

### v4

@@ E.2 Analysis
(Guo-Ren[31]) Under AHPODH, if any attack succeeds in violating a transaction’s safety, then the private mining attack also succeeds in violating that transaction’s safety.
@@
Let $\lambda_{B}=\frac{1}{T_{B}}$ be the block generation rate where $T_{B}$ denotes the block interarrival time. Let $\lambda_{P}$ denote the proof generation rate. Let $\beta$ denote the fraction of honest hashpower in the system. As honest miners need $L$ sequential proofs in the rigged model, we assume
@@
Let $S_{L}\sim NegBin(L,\beta)$ denote the sum of $L$ i.i.d. geometric random variables with success probability $\beta$. Let $S_{\delta_{L}}\sim\sum_{i=1}^{L}Poisson(\lambda_{P}\delta)$ denote the sum of $L$ i.i.d. Poisson random variables with rate $\lambda_{P}\delta$, whereas $S_{\Delta_{2}}\sim\sum_{i=1}^{2}Poisson(\lambda_{P}\Delta)$.
@@
Again, for simplicity, we assume the same rigged model of [31], even though it is not possible for the adversary to mount a balanced attack after $\tau_{c}$ anymore [41]. Making the same mild assumption of ignoring the ledger delays after block $h+2$ for simplicity and incorporating all the proofs to be mined after $\tau_{c}$ into a single proof tree, the number of proofs that the adversary can make up for its deficit is denoted as $M_{deficit}\sim M_{lead}$; see [31, 32, 42, 41].
@@ Appendix F MDP Models for the Proposed Protocol
In this section, we assume the reader is familiar with the $\epsilon$-optimal selfish mining strategies of [5]. We avoid presenting the same average reward ratio MDP model and the results. We simply refer the reader to the relevant parts of the paper [5, Section 2, Section 4.1].
@@
We consider the following modified objective function representing the relative payoff,
@@
Figure 10: Reward split MDP vs Nakamoto MDP [5].
F.1 Simulations
To evaluate the performance of the reward split MDP and compare it with the Nakamoto consensus MDP of [5], we run the MDP models for $\gamma=0,0.2,0.4,0.5,0.6,0.8$ and $\alpha$ values range from $0.02$ to $0.48$ with $0.02$ spacing between them with $10^{-5}$ precision. From Figure 10, it is clear that, reward splitting helps adversary when $\gamma$ is low as matching always gives a half reward when $(a,h,fork)=(1,1,relevant)$, whereas the adversarial relative reward in Nakamoto consensus increases with increasing $\gamma$ since matching succeeds more often. Thus, even when $\gamma=0.8$, the profitability threshold in our protocol stays around $\alpha\approx 0.28$.
@@ Appendix G Transaction Split Rule
It is in the interest of miner $M$ to create an initiator proof referring to its own proposed ledger instead of the ledger offering the most fees if
@@
    If it is successful to mine an initiator referring to the ledger offering $F_{M}$ fees before others, which happens with probability $\alpha$, it shares the initiator immediately. Other miners accept the initiator of $M$ and mine on top of it since it is the chain with the heaviest work.
@@
For the rest of the L1L-1 incremental proofs after the initiator, the miner $M$ creates $(L-1)\alpha$ more proofs on average. The expected transaction fees rewards is
$F_{M}\cdot\alpha\left(r+\frac{1-r}{L}+\frac{1-r}{L}(L-1)\alpha\right)+F_{M^{\prime}}\cdot(1-\alpha)\left(\frac{1-r}{L}(L-1)\alpha\right).$ (27)
On the other hand if it simply tries to mine an initiator proof referring to the ledger offering $F_{M^{\prime}}$ fees, the expected transaction fees rewards is